Calculate the amount of energy needed to convert 11.3 gof water at 35.0 °C to steam at 100.0°C. kJ
Answer #1

Ti = 35.0 oC

Tf = 100.0 oC

here

Cl = 4.184 J/goC

Heat required to convert liquid from 35.0 oC to 100.0 oC

Q1 = m*Cl*(Tf-Ti)

= 11.3 g * 4.184 J/goC *(100-35) oC

= 3073.148 J

Lv = 2260.0 J/g

Heat required to convert liquid to gas at 100.0 oC

Q2 = m*Lv

= 11.3g *2260.0 J/g

= 25538 J

Total heat required = Q1 + Q2

= 3073.148 J + 25538 J

= 28611 J

= 28.6 KJ

Answer: 28.6 KJ